(i) Derive (algebraically) the savings function if the consumption

function is C = 200 + 0.6Y. [4 marks]

II. Graphically plot the savings function derived above. [4 Marks]

Note: In your Graph Indicate the following positions: Y>C; Y<C;

Y=C


1
Expert's answer
2021-05-16T17:42:30-0400

Solution:

i). The saving function with respect to disposable income (Y) is as follows:

S = Y – C

Where:

S = -a + (1 – b) Y

Where: -a = autonomous consumption.

            (1 – b) = marginal propensity to save (MPS).

                  b = marginal propensity to consume (MPC)

             Y = income

S = -a + (1 – b) Y

S = -200 + (1 – 0.6) Y

S = -200 + 0.4Y

Saving function = -200 + 0.4Y


It is depicted as follows on the below graph: